2, what is the total rolling resistance for a dirt haul road going up a 3% grade that fees under wheel load allowing 2 inches of tire penetration?
Answer #1

Solution:-

Given

grade = 3%

Tire penetration = 2 inches

Rolling resistance factor = 40 + 30 * (penetration in inches)

= 40 + 30 *(2)

= 100 lb/ton

Let weight of vehicle = P tons

Rolling resistance of vehicle = Rolling resistance factor * weight of vehicle

= 100 P ( lb )

Grade resistance factor = 20 * grade(%)

= 20 * 3

= 60 lb/ton

Grade resistance = grade resistance facor * weight of vehicle

= 60 P (lb)

Total rolling resistance = 100 P + 60 P = 160 P ( lb ) Abswer